# Fundamental Oscillations of Massive Boson Stars and Distinguishability

### v2

@@ VI Summary
In this work, we focussed on Boson stars, a class of dark stars made of scalar bosons. These are modeled by a scalar field $\phi$ that interacts gravitationally and has self-interaction given by the potential $V(\phi)$. We focus on the simple case of quartic potential given by $V(\phi)=\frac{1}{2}m^{2}\phi^{2}+\frac{1}{4}\lambda\phi^{4}$. The strong-interaction limit ($\Lambda=\lambda M_{Pl}^{2}/4\pi m^{2}/\gg 1$) leads to the formation of a “massive star” whose mass scales as $\propto\sqrt{\lambda}M_{Pl}^{3}/m^{2}$ resembling massive fermionic stars. Such stars are compact objects and can act as BH mimickers, forming a fraction of the DM in the universe.
@@
We evaluate the $f$-mode characteristics for massive BSs within a full general-relativistic setup. We derive the $f$-mode frequency and damping time scaling relations and demonstrate for the first time that the scaled dimensionless $f$-mode characteristics ($f^{\prime}$ and $\tau^{\prime}$) are independent of the model parameters. We report the relations $f^{\prime}-M^{\prime}$, $f^{\prime}-C^{\prime}$, and $\tau^{\prime}-M^{\prime}$, $\tau^{\prime}-C^{\prime}$, which are unique curves for massive BSs. The scaling relations $f=f^{\prime}/(xM_{Pl})$ and $\tau=\tau^{\prime}(xM_{Pl})$ can be used to evaluate the $f$-mode characteristics for any parameter values in the available parameter space. $f^{\prime}$ is an increasing function of $M^{\prime}/C^{\prime}$ whereas $\tau^{\prime}$ is a decreasing one. Thus, the configuration corresponding to the maximum $M^{\prime}=0.06$ or $C^{\prime}=0.16$ was used to derive the maximum value of $f^{\prime}$ given by $f^{\prime}_{max}=0.21$ and a minimum value of $\tau^{\prime}$ given by $\tau^{\prime}_{min}=1900$. This also results in a minimum number of BS $f$-mode oscillations given by $N_{min}=406$.
@@
We compare the BS $f$-mode frequencies with those of NSs and BH $l=2$ QNMs. For the maximum compactness configuration, the frequency is the highest and only depends on the mass, as in the case of BH. For such configurations, the BS $f$-mode frequency is lower than the corresponding BH frequency by a constant factor given by $\mathcal{R_{BS/BH}}=0.22$. Thus, observing BH mimickers with $f$-mode frequency consistently lower than 0.22 times that of the expected BH frequency would be a clear signature of massive BS. We also compare the BS $f$-mode frequency with the corresponding ISCO frequency and find that the ratio of $f$-mode frequency to that of ISCO is larger for BS than NS and BH, given their low compactness. Thus, we expect the effect of the dynamical tides in binary systems of massive BSs to be less prominent.
@@
Flores et al. [50] is the first work to explore non-radial fundamental modes for massive BSs in the strong-coupling limit, followed by Celato et al. [14]. These works explored BS configurations for select model parameters. The scalings with $x$ introduced in this work can be used to compute $f$-mode characteristics for any model parameters using a single solution, establishing that there is no need to study $f$-modes separately for different parameter values.
A list of observational signatures for BSs and a comparison of $f_{ISCO}$ between ECOs and BHs was provided in [21]. This work extends the list to include the fundamental QNM frequencies and demonstrates that $f$-modes can be used as a tool to distinguish BSs from NSs and BHs.
Future: The results presented here can be used for searches of massive BS in the strong-interaction limit where the effect of $f$-modes is important. These would be relevant in the future when $f$-mode oscillations become detectable. We will discuss the detectability of $f$-modes from BSs in the other work [43].
@@
In this work, we have compared BS $f$-modes with those of NSs and BHs. This can be further extended to include NSs with exotic matter such as hybrid stars, strange stars, DM admixed NSs, etc. Here we considered scalar DM with $\phi^{4}$ potential resulting in $C_{max}\approx 0.16$. Higher $C_{max}$ can be achieved with $\phi^{n}$ potential for $n>4$ [35] possibly resulting in a higher $f^{\prime}_{max}$ and a higher $\mathcal{R}$. This can lead to configurations filling the gap between NSs and BHs (see Fig. 3), making it distinguishable from NSs as well. A complete analysis of ECOs would also require consideration of other scalar DM potentials, vector DM, and other ECOs such as fermion stars, gravastars, wormholes, fuzzballs, etc. [13]. The work can also be extended to include spinning BSs.
Acknowledgements.
@@
L.S. and J.S.B. acknowledge support by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) through the CRC-TR 211 ‘Stronginteraction matter under extreme conditions” – project no. 315477589 – TRR 211.
Supplementary material
@@ Appendix A Scaling of $f$-mode equations: Cowling Formalism
We will now see how the $f$-mode equations and the $f$-mode parameters get modified when scaling and scaled EoS are applied. We first look at the equations in Cowling approximation where metric perturbations are neglected, assuming a weak gravitational field. We do not use the Cowling formalism in this work as we solve the equations in full general relativity. However, we demonstrate that the scaling holds for solutions in Cowling approximation as well for the sake of completeness and the benefit of any future works.
@@
Appendix B $f$-mode equations: Full General Relativity
Here, we present the equations used to solve the complex QNM frequencies using full-general relativity, where metric perturbations are coupled to matter. Then we provide the scaling of various quantities that make the equations independent of the model parameters.
@@ B.1 Perturbations Inside the Star
[…]
@@
$\frac{dH_{1}}{dr}$ $=$ $\frac{-1}{r}\left[l+1+\frac{2m}{r}e^{2\lambda}+4\pi r^{2}e^{2\lambda}G\left(p-\epsilon\right)\right]H_{1}+\frac{1}{r}e^{2\lambda}\left[H+K+16\pi G\left(p+\epsilon\right)V\right]\ ,$ (39)
@@
$\left[1-\frac{3m}{r}-\frac{l(l+1)}{2}-4\pi r^{2}Gp\right]H-8\pi r^{2}e^{-\Phi}X+r^{2}e^{-2\lambda}\left[\omega^{2}e^{-2\Phi}-\frac{l(l+1)}{2r}\frac{d\Phi}{dr}\right]H_{1}$ (44)
@@
While solving these equations, we have to impose boundary conditions, i.e., the perturbation functions are finite throughout the interior of the star (particularly at the center, i.e., at $r=0$), and the perturbed pressure ($\Delta p$) vanishes at the surface. Function values at the center of the star can be found using the Taylor series expansion method described in Appendix B of [31] (see also Appendix A of [44]). The vanishing perturbed pressure at the stellar surface is equivalent to the condition $X(R)=0$ (as, $\Delta p=-r^{l}e^{-\Phi}X$). We follow the procedure described in [31] to find the unique solution for a given value of $l$ and $\omega$ satisfying all the boundary conditions inside the star.
@@ B.3 Scaling
- $\zeta^{\phi}=\zeta^{\phi\prime}$
Applying these transformations and using $G=1/M_{Pl}^{2}$ in all the above equations, we find that they can be written completely in terms of the primed quantities, and all the factors of $x$ cancel out. Initially, the $x$ dependence was via the equation of state $p$ and $\epsilon$. Now $p^{\prime}$ and $\epsilon^{\prime}$ do not depend on $x$. Thus, the equations are completely free of the model parameters. For each central density $\rho^{\prime}_{c}$ or equivalently mass $M^{\prime}$, we get a unique solution for $\omega^{\prime}$. The scaling relation $\omega=\omega^{\prime}/(xM_{Pl})$ can then be used to obtain the $f$-mode frequency and damping time for any BS with DM model parameters $(\lambda,m)$, i.e., $\omega(x,M)=\omega^{\prime}(M^{\prime}=M/(xM_{Pl}^{3}))$. $\omega\propto f+i/\tau$. Thus we casn write $f=f^{\prime}/(xM_{Pl})$ and $\tau=\tau^{\prime}(xM_{Pl})$.
@@
Furthermore, the damping time should follow (see, Misner, Thorne, and Wheeler (1973) chapter 36.2); $\tau\sim R^{4}/(GM)^{3}$, which is also consistent with the derived scaling for $\tau$. The damping time is defined to be the inverse of the imaginary part of the complex eigenfrequency, so it should scale as the inverse of $f$. Moreover, the mass-scaled damping time ($GM/\tau$) shows universality with stellar compactness $C=M/R$ [47]. Hence, the initial guess would be that $\tau$ can be scaled as $r$.

### v3

@@ 6 Summary
We have proposed new scaling relations connecting global characteristics of fundamental modes in Boson stars (BS) with microscopic dark matter (DM) parameters. We considered a class of dark stars called Boson stars, modeled by a complex scalar field $\phi$ that interacts gravitationally and has self-interaction given by the potential $V(\phi)$. We focussed on the simple case of quartic potential given by $V(\phi)=\frac{1}{2}m^{2}\phi^{2}+\frac{1}{4}\lambda\phi^{4}$. The strong-interaction limit ($\Lambda=\lambda M_{Pl}^{2}/4\pi m^{2}/\gg 1$) leads to the formation of a “massive star” whose mass scales as $\propto\sqrt{\lambda}M_{Pl}^{3}/m^{2}$, resembling massive fermionic stars. Such stars are compact objects and can act as BH mimickers, forming a fraction of the DM in the universe.
The main highlights of this work:
  • 1.
@@
    We evaluated the $f$-mode characteristics for massive BSs within a full general-relativistic setup. We derived the $f$-mode frequency and damping time scaling relations and demonstrated for the first time that the scaled dimensionless $f$-mode characteristics ($f^{\prime}$ and $\tau^{\prime}$) are independent of the model parameters.
  • 2.
    We report the relations $f^{\prime}-M^{\prime}$, $f^{\prime}-C^{\prime}$, and $\tau^{\prime}-M^{\prime}$, $\tau^{\prime}-C^{\prime}$, which are unique curves for massive BSs. The scaling relations $f=f^{\prime}/(xM_{Pl})$ and $\tau=\tau^{\prime}(xM_{Pl})$ can be used to evaluate the $f$-mode characteristics for any parameter values in the available parameter space.
  • 3.
    $f^{\prime}$ is an increasing function of $M^{\prime}/C^{\prime}$ whereas $\tau^{\prime}$ is a decreasing one. Thus, the configuration corresponding to the maximum $M^{\prime}=0.062$ or $C^{\prime}=0.16$ was used to derive the maximum value of $f^{\prime}$ given by $f^{\prime}_{max}=0.21$ and a minimum value of $\tau^{\prime}$ given by $\tau^{\prime}_{min}=1900$. This also results in a minimum number of BS $f$-mode oscillations given by $N_{min}=406$.
  • 4.
    The derivation of the scaling relations allows the computation of BS QNMs for any parameters and enables a systematic comparison of $f$-mode parameters with other compact objects. We compared the BS $f$-mode frequencies with those of NSs and BH $l=2$ QNMs. We derived the expression for BS $f$-mode frequency that depends solely on the mass and compactness.
  • 5.
@@
    For the maximum compactness configuration, the frequency is the highest and only depends on the mass, as in the case of BHs. For such configurations, the BS $f$-mode frequency was shown to be lower than the corresponding BH frequency by a constant factor given by $\mathcal{R_{BS/BH}}=0.22$. Thus, observing BH mimickers with $f$-mode frequency consistently lower than 0.22 times that of the expected BH frequency would be a clear signature of massive BS.
  • 6.
    We also compared the BS $f$-mode frequency with the corresponding ISCO frequency and found that the ratio of $f$-mode frequency to that of ISCO is larger for BSs compared to BHs, given their low compactness.
@@
[54] is the first work to explore non-radial fundamental modes for massive BSs in the strong-coupling limit, followed by [13]. These works explored BS configurations for select model parameters. The scalings with $x$ introduced in this work can be used to compute $f$-mode characteristics for any model parameters using a single solution, establishing that there is no need to study $f$-modes separately for different parameter values. A list of observational signatures for BSs and a comparison of $f_{ISCO}$ between ECOs and BHs was provided in [19]. This work extends the list to include the fundamental QNM frequencies as a tool to distinguish BSs from NSs and BHs.
The results presented here can be used for searches of massive BS in the strong-interaction limit where the effects of $f$-mode oscillations are important. These would be relevant in the future when $f$-mode oscillations become detectable. With upcoming GW detectors across various frequency bands like LISA (0.1 mHz-1 Hz), LIGO (50 Hz-1000 Hz), and NEMO (2 kHz-4 kHz), a wide range of BS masses could be probed. As the $f$-mode frequency depends on compactness and mass, we get that for $C\gtrsim 0.01$, BSs roughly falling in the mass range of $10^{2}-10^{7}M_{\odot}$ would be observable by LISA, $0.1-10M_{\odot}$ by LIGO and $0.01-1M_{\odot}$ by NEMO. The next-generation detectors CE and ET roughly encompass the range probed by LIGO and NEMO. Although tidal deformability remains the dominant distinguishing parameter in a binary inspiral [46], as it was shown that incorporating higher order effects like the spin-induced quadrupole moment improves the distinguishability [28, 29] and the constraining power of BS microscopic parameters [36, 52], including the effect of $f$-mode oscillations has the potential to improve it further. Distinguishing from NSs in binary remains a challenge, as BSs follow the same $f$-Love universal relations. However, the results are particularly useful as a distinguishability tool in the context of post-merger GW spectroscopy and in future detections of GW bursts from oscillations of isolated systems. Further analyses on the detectability of these $f$-modes are essential. A detailed study of their detectability from isolated GW bursts has been carried out in a separate work [47].
@@
In this work, we have compared BS $f$-modes with those of NSs and BHs. This can be further extended to include NSs with exotic matter such as hybrid stars, strange stars, DM admixed NSs, etc. Here we considered scalar DM with $\phi^{4}$ potential resulting in $C_{max}\approx 0.16$. Higher $C_{max}$ can be achieved with $\phi^{n}$ potential for $n>4$ [37] possibly resulting in a higher $f^{\prime}_{max}$ and a higher $\mathcal{R}$. This can lead to configurations filling the gap between NSs and BHs (see Fig. 3), making it distinguishable from NSs as well. A complete analysis of ECOs would also require consideration of other scalar DM potentials, vector DM, and other ECOs such as fermion stars, gravastars, wormholes, fuzzballs, etc. [12].
Acknowledgements
@@
L.S. and J.S.B. acknowledge support by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) through the CRC-TR 211 ‘Strong interaction matter under extreme conditions’ – project no. 315477589 – TRR 211.
@@ Appendix A Scaling of $f$-mode equations: Cowling Formalism
We will now see how the $f$-mode equations and the $f$-mode parameters get modified when scaling and scaled EoS are applied. We first look at the equations in the Cowling approximation, where metric perturbations are neglected, assuming a weak gravitational field. We do not use the Cowling formalism in this work as we solve the equations in full general relativity. However, we demonstrate that the scaling holds for solutions in Cowling approximation as well, for the sake of completeness and the benefit of any future works.
@@
Appendix B Scaling of $f$-mode equations: Full General Relativity
Here, we present the equations used to solve the complex QNM frequencies using full general relativity, where metric perturbations are coupled to matter. Then we provide the scaling of various quantities that make the equations independent of the model parameters.
@@ B.1 Perturbations Inside the Star
The Lagrangian displacement vector $\textbf{$\zeta$}=(\zeta^{r},\zeta^{\theta},\zeta^{\phi})$ for the polar perturbations of the fluid is given by,
@@
$\frac{dH_{1}}{dr}$ $=$ $\frac{-1}{r}\left[l+1+\frac{2m}{r}e^{2\Lambda}+4\pi r^{2}e^{2\Lambda}G\left(p-\epsilon\right)\right]H_{1}$ (33)
$+\frac{1}{r}e^{2\Lambda}\left[H+K+16\pi G\left(p+\epsilon\right)V\right]\ ,$
@@
$\left[1-\frac{3m}{r}-\frac{l(l+1)}{2}-4\pi r^{2}Gp\right]H-8\pi r^{2}e^{-\Phi}X$ (38)
$+$ $r^{2}e^{-2\Lambda}\left[\omega^{2}e^{-2\Phi}-\frac{l(l+1)}{2r}\frac{d\Phi}{dr}\right]H_{1}$
@@
While solving these equations, we have to impose boundary conditions, i.e., the perturbation functions are finite throughout the interior of the star (particularly at the center, i.e., at $r=0$), and the perturbed pressure ($\Delta p$) vanishes at the surface. Function values at the center of the star can be found using the Taylor series expansion method described in Appendix B of [31] (see also Appendix A of [48]). The vanishing perturbed pressure at the stellar surface is equivalent to the condition $X(R)=0$ (as $\Delta p=-r^{l}e^{-\Phi}X$). We follow the procedure described in [31] to find the unique solution for a given value of $l$ and $\omega$ satisfying all the boundary conditions inside the star.
@@ B.3 Scaling
- $\zeta^{\phi}=\zeta^{\phi\prime}$ -
Applying these transformations and using $G=1/M_{Pl}^{2}$ in all the above equations, we find that they can be written completely in terms of the primed quantities, and all the factors of $x$ cancel out. Initially, the $x$ dependence was via the equation of state $p$ and $\epsilon$. Now $p^{\prime}$ and $\epsilon^{\prime}$ do not depend on $x$. Thus, the equations are completely free of the model parameters. For each central density $\rho^{\prime}_{c}$ or, equivalently, mass $M^{\prime}$, we get a unique solution for $\omega^{\prime}$. The scaling relation $\omega=\omega^{\prime}/(xM_{Pl})$ can then be used to obtain the $f$-mode frequency and damping time for any BS with DM model parameters $(\lambda,m)$, i.e., $\omega(x,M)=\omega^{\prime}(M^{\prime}=M/(xM_{Pl}^{3}))$. $\omega\propto f+i/\tau$. Thus we can write $f=f^{\prime}/(xM_{Pl})$ and $\tau=\tau^{\prime}(xM_{Pl})$.
@@
Furthermore, the damping time should follow (chapter 36.2 of [34]) $\tau\sim R^{4}/(GM)^{3}$, which is also consistent with the derived scaling for $\tau$. The damping time is defined to be the inverse of the imaginary part of the complex eigenfrequency, so it should scale as the inverse of $f$. Moreover, the mass-scaled damping time ($GM/\tau$) shows universality with stellar compactness $C=M/R$ [51]. Hence, the initial guess would be that $\tau$ can be scaled as $r$.